OpenStudy (anonymous):

the electric field strength 5cm from a very long charged wire is 2000N/C. what is the electric field strength 10cm from the wire?

OpenStudy (shayaan_mustafa):

Use equation, E=F/q

OpenStudy (anonymous):

what i have is \[E=1/(4\pi \epsilon _{0}) * Q/r^{2}\]

OpenStudy (shayaan_mustafa):

but you don't have charge. so you can't use this relation.

OpenStudy (anonymous):

\[Q=L \lambda\] and the prof said use 2 for L. not sure why, maybe 10cm/5cm but why do that?

OpenStudy (anonymous):

after i solve for 5cm, i use the value i got for charge density to plug in for 10cm and got 500 N/C. i don't understand why L=2 and if i still use L=2 to solve for 10cm

OpenStudy (shayaan_mustafa):

there are two equations, 1) E=kq/r² find charge from above equation and put it in, 2) E=F/q find Force from above relation

OpenStudy (shayaan_mustafa):

did you got?

OpenStudy (shayaan_mustafa):

why L=2?? may be due to doubling the length.

OpenStudy (anonymous):

not really, i can't recall how to solve for force. force of field on wire or vice versa?

OpenStudy (anonymous):

or are you saying solve for force after i find E?

OpenStudy (shayaan_mustafa):

ok follow me.

OpenStudy (anonymous):

alright

OpenStudy (shayaan_mustafa):

use 1 equation and find Q by substituting values, k=9*10^9 E=2000 r=0.05m

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Q=5.56*10^-10

OpenStudy (shayaan_mustafa):

ok good. Now use equation 2 and find F. by substituting values of, E=2000 q=5.56*10^-10

OpenStudy (anonymous):

should be 1.11*10^-6

OpenStudy (anonymous):

i will try to work more on it you will get a notification if i add more to this later, gotta go, thank you

OpenStudy (shayaan_mustafa):

you are welcome.

OpenStudy (jamesj):

The electric field about a long, straight charged wire is given by \[ E(r) = \frac{\lambda}{2\pi \epsilon_0 r} \] where \( \lambda \) is the charge density. (See for example here: http://farside.ph.utexas.edu/teaching/302l/lectures/node26.html ) Given that, and what you are told about \( E(r) \) for \( r = 0.05 m \), you can then find \( E(r) \) for \( r = 0.10 m \).

OpenStudy (shayaan_mustafa):

hey jamesj who is this person who is reviewing 3rd question of emcrazy14?

OpenStudy (anonymous):

i am still confused. i solved for the charge density given 2000N/C for 5cm but i used 2 for L. L isn't given so i really don't understand why 2. well, it's the same wire the question is asking about so i used it and my answer is 500N/C which seems reasonable...

OpenStudy (jamesj):

From the formula I wrote above for the electric field, you see that it scales as 1/r. So at twice the distance away it should be 1/2, not 1/4

OpenStudy (anonymous):

thx, i had my r squared, should be able to get it now